# Opportunities and Challenges of Generative-AI in Finance

### v1

@@ IV Gen-AI Methodologies In Finance Ecosystem
In this section, we present various ways in which Gen-AI techniques can be trained or fine-tuned for potential use-cases.
IV-A Out-of-Box LLM
@@
In this technique, the users leverage either open source models or Gen-AI/LLM service providers like OpenAI [24], Gemini [28], Microsoft [30] and Perplexity [29] for their tasks. This method requires no training data, and either minimal to no compute resources. If users leverage Gen-AI techniques from service providers, then they do not need to go through the process of deploying Gen-AI models either in house or on cloud.
@@
Further, users can make use of prompt engineering [25] techniques to complete their task. Broadly, there are 2 main varieties of prompt engineering techniques i.e. zero-shot [26] and few-shot [27, 64]. In the case of zero-shot learning, the LLM’s are able to perform the task at hand without seeing any prior examples, but just by leveraging the knowledge present in them. On the other hand in case of few-shot learning, users provide a few examples as part of the prompt for the LLM’s to learn and perform the task at the end of the prompt. In particular, in this technique, the LLM’s leverage their induction heads to perform in-context learning and arrive at the result[118, 117]. In summary, the goal is to nudge the state of LLM into the right direction which provides the best response for the given task at hand.
Prompt engineering techniques for task completion is great for users with minimal data and/or compute resources. However, there are a couple of drawbacks involved in using this method.
  1. 1.
    For use cases where the LLM’s are unable to learn from in-context examples (usually happens in the case of using smaller LLM’s), then it is better to fine-tune the model for the specific task.
  2. 2.
    Including examples for the LLM’s to learn from takes up the space context length to provide task relevant prompts.
Both the drawbacks are now becoming irrelevant because of quantized models (see below) and almost infinite context length (1M+ context length in Gemini [28]).
@@ IV-B Fine-tuning
In scenarios, where utilizing LLM’s out of the box is not working for the task at hand, then users could leverage fine-tuning to train the LLMs to perform better at the specific tasks the user has in mind. At a high level, fine-tuning techniques can be divided into the following categories.
@@ IV-B2 Task specific Fine-tuning
Task specific fine-tuning [32] is prone to exhibit a phenomenon called catastrophic forgetting. In catastrophic forgetting, the underlying LLM has totally forgotten the knowledge of the world it had obtained as part of its pre-training and its performance on the other tasks after task specific fine-tuning is much worse than its performance on the same tasks before fine-tuning. In order to circumvent the issue of catastrophic forgetting, users can employ multi-task instruction fine-tuning or employ parameter efficient fine-tuning described in section IV-B3.
@@ IV-B3 Parameter Efficient Fine-tuning
Instruction Fine-Tuning IV-B1 or Task specific fine-tuning IV-B2 are resource intensive and are plagued with catastrophic forgetting. In order to circumvent both these issues users can leverage Parameter Efficient Fine-tuning (PEFT) [33]. Two most commonly used techniques in PEFT are LoRA [6] which rely on re-parameterization technique and soft prompts which add additional trainable layers to the LLM. We will now describe both of them in detail.
@@
    LoRA introduces a new way to train the LLM’s by retaining their pre-training knowledge. The weights of the LLM are frozen and new low rank decomposition matrices are added to every layer in the transformer. The authors of LoRA have achieved good results with having the low-rank of the decomposition matrices as low as 1 or 2. Since these additional matrices are very small in size (order of MBs), multiple LoRA models can be trained and stored. Since these low rank decomposition matrices are quite small in comparison to the original LLM weights, and as part of training no optimizer states are stored for the LLM weights, very minimal compute resources are required during training of LoRA models.
@@
    Soft Prompts[65] fall under the additive method paradigm where no weights of the model are changed. Instead of modifying the weights of the pre-trained LLM’s, soft prompts rely on prepending trainable tokens or soft prompts to the input tokens during training for a given task. The loss is propagated all the way to these trainable tokens and an efficient representation is learnt for the task at hand. The idea of these soft tokens is to choose the right vectors for a given space on the N-dimensional hypersphere of the embedding vector space. Since these soft tokens are very light-weight, multiple such soft tokens can be learnt similar to LoRA.
@@ IV-C Agentic Systems
Even though LLM’s have made great strides in their ability to understand natural languages (via zero-shot or few shot examples), LLM’s still struggle in providing accurate up to date information, hallucinate answers[34, 35] or are unable to perform precise mathematical calculations[59]. A very simple solution is to provide the ability for LLM’s to utilize external tools such as search engine, calculators etc which help the LLM’s in overcoming their major drawbacks.
Toolformer[60] is one of the examples for LLM’s that utilize tools to overcome the above mentioned drawbacks. In the case of Toolformer, the LLM training involves three steps
  1. 1.
    Given a dataset C = $\{x^{1},x^{2},...,x^{N}\}$ of plain text, utilizing the in-context learning ability of the already trained LLM, each of the plain text is augmented with potential API(tool) calls.
  2. 2.
    API/tool calls are executed and their responses are collected. Responses which do not reduce the loss (wrong API/tool call responses) are discarded and and the rest of the responses are stored to be used as dataset for fine-tuning the LLM for tool usage.
  3. 3.
    Utilizing this newly augmented dataset the LLM is then fine-tuned.
@@ IV-D Quantization
Traditionally models were trained and deployed using 32-bit floating point numbers to represent the parameters of the model. With the exponential increase in the number of parameters in LLM’s, one major concern that arises is the amount of time taken for inference. In order to speed up the inference time researchers leverage quantization to represent the parameters of the model using less bits (float16, bfloat16 or int8) without much loss in accuracy. Users utilize one of the two quantization schemes
@@ V Gen-AI Applications in finance ecosystem
Gen-AI, is still in its nascent stages and implementation within finance sector comes with it’s own set of risks and challenges. However, the trans-formative nature of this technology is driving many applications in this domain. In this section, we will highlight such applications both in the industry and the research community.
@@ V-A Numerical Reasoning
In recent times, datasets with complex numerical reasoning tasks have been proposed. In particular, FinQA [75] is based on the earnings reports of S&P 500 companies. The questions in FinQA, such as “Considering the weighted average fair value of options, what was the change of shares vested from 2005 to 2006?” and “What was the net change in tax positions in 2014?”, require information from both tables and unstructured texts to answer. Similarly, ConvFinQA [76], an extension of FinQA, is a multi-turn conversational question-answering dataset over financial reports, consisting of 3,892 conversations with 14,115 questions.
GPT-4 achieved an accuracy of 78% on FinQA and 76% on ConvFinQA dataset which is much higher than the average human [86]. This ability has enabled real world used cases like Fintool [79] - an AI equity research tool that is engineered to discover financial insights about companies.
Credit Karma by Intuit [85], is making use of the Intuit Assist to launch Gen-AI experiences which include asking questions about a user’s personal finance, understanding spend and personal financial roadmap, save money by using offers and much more. Mastercard reports that their new transaction decision technology [82] based on LLM’s could help financial institutions improve their fraud detection rates by 20%, on average.
In RiskLabs, Cao et al. [45] explored a new framework that leverages Gen-AI to analyze and predict financial risks. It uniquely combines different types of financial data, including textual and vocal information from Earnings Conference Calls (ECCs), market-related time series data, and contextual news data and showcases the critical role of LLM’s in financial risk assessment and opens new avenues for their application in this field.
Capturing multi-modal input via Gen-AI is gaining traction in the finance domain. Chang et al [80] proposed usage of key technologies of Gen-AI into financial market prediction which can effectively capture the complexity of financial market data, and with a high degree of accuracy. Lezhi et al [81] showcase how Multimodal Gen-AI could be used for fundamental investment research. Through fine-tuning methods applied to a base model (Llama2), they develop an AI agent that can assist investors in tasks such as understanding market conditions, generating investment ideas, and formatting results with stock recommendations. JP Morgan share their LLM called DocLLM, which is a layout-aware generative language model for multimodal document understanding which can include forms, invoices, receipts, reports, contracts, and other similar records, often carry rich semantics at the intersection of textual and spatial modalities. [84]
@@ V-B Trading
Gen-AI has also seen applications in trading yielding impressive results in research settings.
In particular, Wu et al. (2023) introduced BloombergGPT [3], a 50-billion parameter large language model designed specifically for the financial domain. This was trained on a combination of an internal dataset built from Bloomberg’s extensive archives called FinPile along with publicly available datasets. They demonstrated the effectiveness of training a model that excels in tasks like sentiment analysis and question answering in the financial context, while maintaining proficiency in general language tasks.
Further, Mai (2024) introduced StockGPT [2], which was trained on numeric data containing around 70 million daily U.S. stock returns over nearly 100 years. The stock return series was used as token sequences helping StockGPT learn predictive patterns through the attention mechanism. The model demonstrates remarkable performance, achieving high annual returns and a Sharpe ratio of 6.5 on a hold-out test sample from 2001 to 2023.
Furthermore, FinBERT [17] is a pre-trained language model specifically for the financial domain. It outperforms generic language models on financial sentiment classification tasks. Financial sentiment classification for trading using various GPT’s has been explored yielding results with sharpe ratio over 3. [120]
Additionally, Lopez et al. [121] showcased that incorporating advanced language models into the investment decision-making process can yield more accurate predictions and enhance the performance of quantitative trading strategies. Fatourosa et al. [122] propose that a portfolio rebalanced monthly using the buy/sell signals generated from an LLM can outperform a passive index by 10-30%.
@@ V-C Summarization and Assistive Applications
Gen-AI techniques have also been applied in text summarization and assistive applications. In particular, Xiao et al.[74] discussed how there was a clear preference among human evaluators for LLM-generated summaries over human-written summaries or summaries generated by fine-tuned models.
Further, JP Morgan uses a generative AI tool that is designed to serve as a ‘research analyst’ for over 50,000 employees, aiding in various tasks that enhance productivity and decision-making within the firm [83]. Models like SaulLM-7B [72] based on mistral 7B architecture, are designed and distributed that help in legal text comprehension and generation.
Companies like Kudos and Max-Rewards understand credit card offerings and combine them with user’s spends to suggest the best credit cards for maximising rewards. Morgan Stanley’s COIN (Contract Intelligence) uses AI to review legal documents and extract relevant information.
V-D Other Applications
Positive customer interactions aided by Gen-AI leads to increased brand satisfaction, trust and engagement [124] which would in-turn help institutions acquire customers and market and sell products better.
Leading CRM companies like Salesforce [69], Hubspot [70], Zendesk [71] and others offer AI based chatbots for better customer experience and reducing support volumes.
Ferrero et al. [125] showcase how GenAI is increasingly applied to customer service operations and provide some brand response strategies to mitigate the challenges with this application. Timothy et al. [77] proposes a conceptual framework that analyzed the unique demands of suggestion contexts and built a prototype called Omnifill that uses LLM’s to generate suggestions for various form-filling tasks.
TurboTax, a tax preparation software, uses Intuit Assist [78], which is a generative AI-powered financial assistant. It uses a combination of the company’s prior tax prep experience, data and documents from the filer, and current tax code to do the hard work for tax filing.
@@ VI Conclusion
Machine Learning and data mining techniques have led to massive impact via automation of repetitive tasks, improved precision and accuracy, enhanced decision making, personalized and tailored customer experience and operational scalability across various domains (e.g. healthcare, finance and automation). However, aforementioned techniques are limited in their ability to understand and reason languages.
The aforementioned challenges were overcome by the arrival of Gen-AI techniques. Gen-AI techniques are able to improve understanding of context and nuances in language modeling, translation between languages, handle large volumes of data, provide fast and low-latency responses and can be fine-tuned for various tasks and domains. In this manuscript we have provided a brief overview of the application of Gen-AI techniques within the finance ecosystem.
In particular, we discussed the various opportunities and challenges associated with the widespread adoption of Gen-AI techniques in the finance domain. We also present on how such large models can be trained and be made available by overcoming challenges stemming from critical issues such as hallucination, data-quality and computational costs. Lastly, we provide application areas where these models are or can be used.
To the best of our knowledge, this is the first work which comprehensively summarizes the usage of Gen-AI techniques in the finance domain. This summarization would not only provide a clear overview of the areas where substantial work has been completed but also pinpoint areas where future efforts can be prioritized, potentially accelerating the development and adoption of Gen-AI in finance. Moreover, this analysis could serve as a valuable bridge between finance and other domains, facilitating the cross-pollination of innovative ideas and best practices, ultimately benefiting a wider range of industries and applications.

### v3

@@ IV Gen-AI Methodologies
This section presents various ways to train or fine-tune Gen-AI techniques for potential use-cases in finance.
IV-A Out-of-Box
@@
In this technique, the users leverage either open source models or Gen-AI service providers like OpenAI, Gemini, Microsoft and Perplexity for their tasks. This method requires no training data, and either minimal to no compute resources.
@@
Further, users can make use of prompt engineering [45] techniques to complete their task. Broadly, there are two main varieties of prompt engineering techniques i.e. zero-shot [46] and few-shot [47]. In the case of zero-shot learning, Gen-AI techniques are able to perform the task at hand without seeing any prior examples, but just by leveraging the knowledge present in them. On the other hand in case of few-shot learning, users provide a few examples as part of the prompt for the Gen-AI techniques to learn and perform the task at the end of the prompt. In particular, in this technique, the Gen-AI techniques can leverage their induction heads [48] to perform in-context learning [49] and arrive at the result.
@@ IV-B Fine-tuning
In scenarios, where utilizing LLMs out of the box is not working for the task at hand, then users could leverage fine-tuning to train the LLMs to perform better at the specific tasks the user has in mind. At a high level, fine-tuning techniques can be divided into the following categories.
@@ IV-B2 Task specific Fine-tuning
Task specific fine-tuning [51] is prone to exhibit a phenomenon called catastrophic forgetting. In catastrophic forgetting, the underlying LLM has forgotten the knowledge of the world it had obtained as part of its pre-training and its performance on the other tasks after task specific fine-tuning is much worse than its performance on the same tasks before fine-tuning. In order to circumvent the issue of catastrophic forgetting, users can employ multi-task instruction fine-tuning or employ parameter efficient fine-tuning described in section IV-B3.
@@ IV-B3 Parameter Efficient Fine-tuning
Instruction Fine-Tuning IV-B1 or Task specific fine-tuning IV-B2 are resource intensive and are plagued with catastrophic forgetting. In order to circumvent both these issues users can leverage Parameter Efficient Fine-tuning (PEFT) [52]. Two most commonly used techniques in PEFT are LoRA [35] which rely on re-parametrization technique and soft prompts which add additional trainable layers to the LLM.
@@
    LoRA introduces a new way to train the LLMs by retaining their pre-training knowledge. The weights of the LLM are frozen and new low rank decomposition matrices are added to every layer in the transformer.
@@
    Soft Prompts [53] fall under the additive method paradigm where no weights of the model are changed. Instead of modifying the weights of the pre-trained LLMs, soft prompts rely on prepending trainable tokens or soft prompts to the input tokens during training for a given task. The loss is propagated all the way to these trainable tokens and an efficient representation is learnt for the task at hand. The idea of these soft tokens is to choose the right vectors for a given space on the N-dimensional hypersphere of the embedding vector space. Since these soft tokens are very light-weight, multiple such soft tokens can be learnt similar to LoRA.
@@ IV-C Agentic Systems
Even though LLMs have made great strides in their ability to understand natural languages (via zero-shot or few shot examples), LLMs still struggle in providing accurate up to date information, hallucinate answers [54, 55] or are unable to perform precise mathematical calculations [56]. A very simple solution is to provide the ability for LLMs to utilize external tools such as search engine, calculators etc which help the LLMs in overcoming their major drawbacks.
@@ IV-D Quantization
Traditionally models were trained and deployed using 32-bit floating point numbers to represent the parameters of the model. With the exponential increase in the number of parameters in LLMs, one major concern that arises is the amount of time taken for inference. In order to speed up the inference time researchers leverage quantization to represent the parameters of the model using less bits (float16, bfloat16 or int8) without much loss in accuracy. Users utilize one of the two quantization schemes
@@ V Gen-AI Applications in finance ecosystem
Gen-AI techniques are still in its nascent stages and its implementation within finance sector comes with its own set of risks and challenges. However, the transformative nature of this technology is driving many applications in this domain. In this section, we will highlight such applications both in the industry and the research community.
@@ V-A Numerical Reasoning
Several datasets consisting of complex numerical reasoning tasks have been proposed to evaluate the efficacy of Gen-AI techniques. For example, FinQA [60], is a dataset with Question-Answer pairs over Financial reports written by financial experts. FinQA, consist of questions such as “Considering the weighted average fair value of options, what was the change of shares vested from 2005 to 2006?”. Similarly, ConvFinQA [61], an extension of FinQA, is a multi-turn conversational question-answering dataset over financial reports, consisting of 3,892 conversations with 14,115 questions. GPT-4 achieved an accuracy of 78% on FinQA and 76% on ConvFinQA dataset which is much higher than the average human [62]. This ability has enabled real world used cases like Fintool [63] - an AI equity research tool that is engineered to discover financial insights about companies.
Credit Karma [64], is making use of the Intuit Assist to launch Gen-AI experiences which include asking questions about a user’s personal finance, understanding spend and personal financial roadmap.
@@ V-B Trading
Gen-AI techniques have also seen applications in trading yielding impressive results in research settings. In particular, Wu et al. (2023) introduced BloombergGPT [28], a 50-billion parameter Gen-AI technique, designed specifically for the financial domain. They demonstrated the effectiveness of their proposed technique for sentiment analysis, financial question answering, while maintaining proficiency in general language tasks.
Mai (2024) proposed StockGPT [65], which uses token sequences to learn predictive patterns via the attention mechanism there by demonstrating substantial impact. Further, Lopez et al. [66] showcased that incorporating advanced Gen-AI techniques into the investment decision-making process can yield accurate predictions and enhance the performance of quantitative trading strategies. Furthermore, Fatourosa et al. [67] proposed that a portfolio rebalanced monthly using the buy/sell signals generated from Gen-AI technique can outperform a passive index by 10-30%.
Additionally, Lezhi et al. [68] showcased how multimodal Gen-AI techniques could be used for fundamental investment research. Through fine-tuning methods applied to a base model (Llama2), they developed an AI (Artificial Intelligence) agent that can assist investors in tasks such as understanding market conditions, generating investment ideas, and formatting results with stock recommendations.
@@ V-C Summarization and Assistive Applications
Gen-AI techniques have also been applied in text summarization and assistive applications. In particular, Xiao et al. [69] observed a clear preference among human evaluators for LLM-generated summaries over human-written summaries. Gen-AI techniques like SaulLM-7B [70] based on mistral 7B architecture, are designed for legal text generation and comprehension.
Further, JP Morgan uses a Gen-AI toolkit, designed to serve as a ‘research analyst’, aiding in various tasks that enhance productivity and decision-making within the firm [71]. Furthermore, companies like Kudos and Max-Rewards understand credit card offerings and combine them with user’s spends to suggest the best credit cards for maximising rewards. Morgan Stanley’s COIN (Contract Intelligence) uses AI to review legal documents and extract relevant information.
Leading CRM companies such has Salesforce [72], Hubspot [73], Zendesk [74] offer AI based chatbots for better customer experience. Further, Intuit has integrated, a Gen-AI powered financial assistant called intuit assist into TurboTax [75], that uses a combination of the company’s prior tax preparation experience, data and documents from the filer, and current tax code for tax filing.
V-D Risk Monitoring
Mastercard [76] discussed how Gen-AI techniques could help financial institutions improve their fraud detection rates by 20%, on average. Further, Cao et al. [77] explored a new framework that leverages Gen-AI to analyze and predict financial risks. It uniquely combines different types of financial data, including textual and vocal information from Earnings Conference Calls (ECCs), market-related time series data, and contextual news data and showcases the critical role of LLMs in financial risk assessment and opens new avenues for their application in this field.
@@ VI Conclusion
Compared to the traditional machine learning and data mining approaches, Gen-AI techniques are able to improve understanding of context and nuances in language modeling, translation between languages, handle large volumes of data, provide fast and low-latency responses and can be fine-tuned for various tasks and domains. In this manuscript we have provided a brief overview of the application of Gen-AI techniques within the finance ecosystem.
Further, we discussed the various opportunities and challenges associated with the widespread adoption of Gen-AI techniques in the finance domain. We also discuss techniques which can be used to develop Gen-AI models by overcoming challenges stemming from critical issues such as hallucination, data-quality and computational costs. Lastly, we provide application areas in Finance where these Gen-AI models are or can be used.
To the best of our knowledge, this is the first work which comprehensively summarizes the usage of Gen-AI techniques in the finance domain. This summarization would not only provide a clear overview of the areas where substantial work has been completed but also pinpoint areas where future efforts can be prioritized, potentially accelerating the development and adoption of Gen-AI techniques in finance. Moreover, this analysis could serve as a valuable bridge between finance and other domains, facilitating the cross-pollination of innovative ideas and best practices, ultimately benefiting a wider range of industries and applications.